OpenStudy (anonymous):

2x^2 times 4x^3/2 times 4x^-2

OpenStudy (jhannybean):

\[\large2x^2 \cdot 4x^{3/2}\cdot 4x^{-2}\]\[\large (2\cdot 4\cdot 4)(x^{(2+(3/2)-2)})\]

OpenStudy (blurbendy):